Find the square of the following numbers. 
71
Find the square of the following numbers. 
71

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(71)^2
= (70+1)^2
= (70)^2+ (1)^2+2×70×1 [Since, (a+b)^2= a^2+b^2+2ab]​
= 4900+1+140
= 5041

​Thanks